OpenStudy (anonymous):

the cubed root of 24a^5

OpenStudy (anonymous):

\[\sqrt[3]{24a ^{5}}= \sqrt[3]{8\times3\times a ^{5}}=\sqrt[3]{8} \times \sqrt[3]{3\times a ^{5}}=2 \times \sqrt[3]{3\times a ^{5}} \]